# An exact dimension-reduced dynamic theory for developable surfaces and curve-fold origami

### v2

@@ 3.2 A nonlinear dynamic bi-rod model for curve-fold origami
which governs the cross-sectional orientation of the two equivalent rods. The Darboux frames are also shown in Fig. 4(a). According to Eq. (44), the angular velocity of the Darboux frame $\mathbf{\Omega}_{\pm}$ relates to the angular velocety of the Frenet frame $\mathbf{\Omega}_{f}$, following
@@
Define $\mathbf{F}_{\Sigma}=\mathbf{F}_{+}+\mathbf{F}_{-}$ and $\mathbf{M}_{\Sigma}=\mathbf{M}_{+}+\mathbf{M}_{-}$. Equation (47) can be simplified into
@@ 4.1 Dynamics of a curve-folded plate
Figure 5: (a) FEM result of the dynamics of a curve-folded plate. (b) Comparison between the theoretical results and FEM results.
A curve-folded plate is a basic cell of curve-fold origami. In this example, we analyse a freely developing curved-folded plate where the fold starts at the middle point along the length. The specific loading process is to press the two long edges to fold the plate statically, and then release the compression to unfold it dynamically. Previous studies indicate that the generators remain parallel in the static deformation (Lee et al., 2018). Here, we first assume that this conclusion remains applicable in dynamics and will verify it subsequently. Under such an assumption, the geometric variants are
@@
where $\eta_{\pm}$ and $v_{0\pm}$ are given in Eq. (50). Solving Eq. (51) yields $\mathbf{\Omega}_{-}\times\mathbf{l-}=0$, indicating that the solution does not violate the assumption that the generators remain parallel and the panels deform cylindrically. Since the dynamic process is uniquely determined by Eq. (47), the solution of Eq. (51) is the unique solution that satisfies Eq. (32) in this example. The process is also simulated in finite element analysis, where the mechanical parameters are selected as $D\kappa_{g}/k_{c}=1$, and the geometric parameters are $L/W=1/2$. The good agreement between the theoretical and numerical results are demonstrated in Fig. 5.
@@ 4.2 Dynamics of single-vertex multi-curve-fold origami
Figure 6: (a) FEM result of the dynamics of the single-vertex curved origami. (b) Comparison between the theoretical results and FEM results.
@@
While the proof is purely geometric and does not introduce any mechanical variables, the conclusion can be extended to dynamical situations. Given the rotation angle at the vertex, the geometry of the folds can be calculated, which is independent of the elasticity according to the result in the static study. Comparison between the FEM results and the theoretical results derived from Eq. (56) is shown in Fig. 6. The good agreement validates the conclusion in dynamical situations.
@@ 4.3 Dynamics of a cut-then-fold ring origami
Figure 7: (a) Finite element result of the dynamics of a cut-then-folded ring. (b) Comparison between the theoretical results and FEM results.
@@
Numerically solving Eq. (57) yields the curvature and torsion distributions at any time $t$. The theory indicates that the torsion of the cut ring gradually decreases while the curvature increases. This behavior is reproduced in finite-element simulations, as shown in Fig. 7, where the mechanical parameters satisfy $D\kappa_{g}/k_{c}=1$ and $\phi_{0}=\pi$. A direct comparison between the theoretical and simulated curvature and torsion is also presented in Fig. 7, which demonstrates good quantitative agreement.
5 Conclusion
@@
We believe that our dynamical theory for developable surfaces and curved origami, including the equivalent rod models, can deepen understanding of the physics of the curve-fold origami system and provide new insights into shape programming. Our theoretical framework can be naturally extended to developable surfaces with reinforcing ribs, such as those in space antennas. Another future direction is to generalize the dynamic framework for developable surfaces that have intersections of generators. For example, a closed Möbius ring cannot avoid the intersection of generators (Starostin and van der Heijden, 2007). We anticipate that our work can contribute to the design of complex dynamic curve-fold origami structures and developable surfaces in the fields of robotics, metamaterials, and architecture.
@@ Appendix A Dynamical equation in the Eulerian framework
where $\mathbf{p}^{*}$ and $\mathbf{p}_{\theta}^{*}$ are the unit-length momentum and angular momentum of the region $(S,S+{\textrm{d}}S)$. The illustration of the region at the time $t$ and $t+{\textrm{d}}t$ is shown in Fig. A.1. Notice that the mass in the region $(S,S+{\textrm{d}}S)$ varies over time. The net momentum and angular momentum input $\dot{\mathbf{p}}_{in}^{*}$ and $\dot{\mathbf{p}}_{in\theta}^{*}$ are considered, which is introduced by the gray region shown in Fig. A.1.
@@ Appendix B Component form of the dynamical equations in the fixed frame
In this appendix, we transform the dynamical equations into component forms expressed in xyzxyz frame, for the convenience of numerical solution.
@@
For a single developable surface, the equations and the process are similar. The specific form of the dynamic equations is similar to Eq. (B.7) by removing the summation and the negative sign, and the fourth equation in Eq. (B.7) no longer exists. Then the variables are $x,y,z,\psi,\theta,\varphi,F_{x},F_{y},F_{z},m_{3}$, and there are 10 equations with 10 variables, which can thus be solved numerically.

### v3

@@ 3.2 A nonlinear dynamic bi-rod model for curve-fold origami
which governs the cross-sectional orientation of the two equivalent rods. The Darboux frames are also shown in Fig. 5(a). According to Eq. (49), the angular velocity of the Darboux frame $\mathbf{\Omega}_{\pm}$ relates to the angular velocity of the Frenet frame $\mathbf{\Omega}_{f}$, following
@@
Define $\mathbf{F}_{\Sigma}=\mathbf{F}_{+}+\mathbf{F}_{-}$ and $\mathbf{M}_{\Sigma}=\mathbf{M}_{+}+\mathbf{M}_{-}$. Eq. (52) can be simplified into
@@ 4.1 Dynamics of a curve-folded plate
A curve-folded plate is a basic cell of curve-fold origami. In this example, we analyse a freely developing curved-folded plate (length: $L$, width: $b$, thickness: $h$) where the fold starts at the middle point along the length. The specific loading process is to press the two long edges to fold the plate statically, and then release the compression to unfold it dynamically. Previous studies indicate that the generators remain parallel in the static deformation (Lee et al., 2018). Here, we first assume that this conclusion remains applicable in dynamics and will verify it subsequently. Under such an assumption, the geometric variants are
@@
where $\eta_{\pm}$ and $v_{0\pm}$ are given in Eq. (55). Solving Eq. (56) yields $\mathbf{\Omega}_{-}\times\mathbf{l-}=0$, indicating that the solution does not violate the assumption that the generators remain parallel and the panels deform cylindrically. Since the dynamic process is uniquely determined by Eq. (52), the solution of Eq. (56) is the unique solution that satisfies Eq. (32) in this example. The process is also simulated in finite element analysis, where the mechanical parameters are selected as $D\kappa_{g}/k_{c}=1$, and the geometric parameters are $L/b=1/2$. The good agreement between the theoretical and numerical results is demonstrated in Fig. 6(b). Fig. 6(c) illustrates the conversion between elastic energy and kinetic energy, which further demonstrates that the elastic energy is completely converted into kinetic energy and the origami structure is fully flattened at time $T$. $E_{0}$ in Fig. 6 is used to nondimensionalize the energy, defined as $E_{0}=0.5D\kappa_{g}^{2}Lb$.
Figure 6: (a) FEM result of the dynamics of a curve-folded plate. (b) Comparison between the theoretical results and FEM results. (c) FEM results of the evolution of kinetic and elastic energy. $T$ is the characteristic time shown in (a). $\kappa_{g}$ is the initial geodesic curvature of the crease, and $E_{0}$ is defined as $E_{0}=0.5D\kappa_{g}^{2}Lb$ to nondimensionalize the energy.
@@ 4.2 Dynamics of single-vertex multi-curve-fold origami
Figure 7: (a) FEM result of the dynamics of the single-vertex curved origami. (b) Comparison between the theoretical results and FEM results. $T$ is the characteristic time between the initial state and the flat state. $\kappa_{g}$ is the geodesic curvature of the crease, which is used to nondimensionalize the curvature.
@@
While the proof is purely geometric and does not introduce any mechanical variables, the conclusion can be extended to dynamical situations. Given the rotation angle at the vertex, the geometry of the folds can be calculated, which is independent of the elasticity according to the result in the static study. The comparison between the FEM results and the theoretical results derived from Eq. (61) is shown in Fig. 7. The good agreement validates the conclusion in dynamical situations.
@@ 4.3 Dynamics of a cut-then-fold ring origami
Figure 8: (a) Finite element result of the dynamics of a cut-then-folded ring. (b) Comparison between the theoretical results and FEM results. $T$ is the characteristic time shown in (a). $\kappa_{g}$ denotes the initial geodesic curvature of the crease, which is used to nondimensionalize the curvature and torsion.
@@
Numerically solving Eq. (62) yields the curvature and torsion distributions at any time $t$. The theory indicates that the torsion of the cut ring gradually decreases while the curvature increases. This behavior is reproduced in finite-element simulations, as shown in Fig. 8, where the mechanical parameters satisfy $D\kappa_{g}/k_{c}=1$ and $\phi_{0}=\pi$. A direct comparison between the theoretically predicted and numerically simulated curvature and torsion is presented in Fig. 8, showing reasonably good quantitative agreement overall. Nevertheless, noticeable discrepancies persist in this example. We attribute this primarily to the numerical difficulty of handling the abrupt geometric boundary change induced by the initial cut. In addition, the assumption underlying Eq. (62) may break down even in the narrow-ribbon limit (Starostin and van der Heijden, 2007).
5 Discussion
@@
We believe that our dynamical theory for developable surfaces and curved origami, including the equivalent rod models, can deepen understanding of the physics of the curve-fold origami system and provide new insights into shape programming. Our theoretical framework can be naturally extended to developable surfaces with reinforcing ribs, such as those in space antennas. For example, we may optimize a developable curved surface as part of an antenna. Within our dynamic framework, any admissible reference curve can be selected to parameterize and describe the motion of a developable surface. By choosing the reinforced rib as the reference curve, the dynamics of the rib can be naturally incorporated into the governing equations of our model, leading to a unified set of dynamic equations for a rib-reinforced developable surface. Since the present dynamic model for developable surfaces and the classical Kirchhoff rod model are both formulated as one-dimensional rod theories, their coupling is conceptually straightforward and mathematically consistent. Another important direction for future work is to extend the present dynamic framework to developable surfaces with intersecting generators and more complex generator distributions. For instance, a closed Möbius ring necessarily involves intersections of generators (Starostin and van der Heijden, 2007), leading to geometric singularities that require a refined treatment. Moreover, dynamic bifurcation phenomena, similar to those reported in (Wen et al., 2025), are expected to emerge in these more intricate configurations. We anticipate that our work can contribute to the design of complex dynamic curve-fold origami structures and developable surfaces in the fields of robotics, metamaterials, and architecture.
@@ Appendix A Dynamical equation in the Eulerian framework
where $\mathbf{p}^{*}$ and $\mathbf{p}_{\theta}^{*}$ are the unit-length momentum and momentum of the region $(S,S+{\textrm{d}}S)$. The illustration of the region at the time $t$ and $t+{\textrm{d}}t$ is shown in Fig. A.1. Notice that the mass in the region $(S,S+{\textrm{d}}S)$ varies over time. The net momentum and angular momentum input $\dot{\mathbf{p}}_{in}^{*}$ and $\dot{\mathbf{p}}_{in\theta}^{*}$ are considered, which is introduced by the gray region shown in Fig. A.1.
@@ Appendix B Component form of the dynamical equations in the fixed frame
In this appendix, we transform the dynamical equations into component forms expressed in the xyzxyz frame, for the convenience of numerical solution. An algorithmic flow description is summarized in the end, and the corresponding equations are recalled.
@@
For a single developable surface, the equations and the process are similar. The specific form of the dynamic equations is similar to Eq. (B.7) by removing the summation and the negative sign, and the fourth equation in Eq. (B.7) no longer exists. Then the variables are $x,y,z,\psi,\theta,\varphi,F_{x},F_{y},F_{z},m_{3}$, and there are 10 equations with 10 variables, which can thus be solved numerically. The process is illustrated in the flowchart below.